OpenStudy (anonymous):

A mountain has a rise of 8,480 feet from its base. If there is a 2 degree drop in temperature for every rise of 530 feet in altitude, how many degrees difference is there between the ground temperature and the temperature at the peak? (Points : 1) 32 1060 16 4240

OpenStudy (anonymous):

@briannamooney

OpenStudy (anonymous):

A is your answer.

OpenStudy (anonymous):

8480/530 = 16 . 16*2 = 32

OpenStudy (anonymous):

(-12)(5) = (Points : 1) 60 -60 -2.4 2.4 Question 15. 15. 25 ÷ (-5) = (Points : 1) 30 -5 -30 5

OpenStudy (anonymous):

Q1) -60

OpenStudy (anonymous):

25 divided by (-5)

OpenStudy (anonymous):

G2)-5

OpenStudy (anonymous):

Medal plz :)

OpenStudy (anonymous):

Simplify: -(9y - 5). (Points : 1) 5 + 9y -9y - 5 5 -9y 9y + 5 @briannamooney

OpenStudy (anonymous):

C